# Beyond Overcorrection: Evaluating Diversity in T2I Models with DivBench

### v1

@@ Abstract
Current diversification strategies for text-to-image (T2I) models often ignore contextual appropriateness, leading to over-diversification where demographic attributes are modified even when explicitly specified in prompts. This paper introduces DivBench, a benchmark and evaluation framework for measuring both under- and over-diversification in T2I generation. Through systematic evaluation of state-of-the-art T2I models, we find that while most models exhibit limited diversity, many diversification approaches overcorrect by inappropriately altering contextually-specified attributes. We demonstrate that context-aware methods, particularly LLM-guided FairDiffusion and prompt rewriting, can already effectively address under-diversity while avoiding over-diversification, achieving a better balance between representation and semantic fidelity. [^1] We publicly release our data and code at huggingface.co/DivBench.
@@ 2.1 DivBench
We describe the overall setup in Fig. 1. We manually collected and annotated samples, divided into two main categories: (i) examples without assigned demographic attributes (e.g., “farmer” or “housekeeper”), which test whether models naturally reflect inclusive and diverse outputs, and (ii) examples with clearly defined demographic constraints (e.g., “actress” or “Maasai person”), which test whether models introduce inappropriate or historically inaccurate diversity, referred to as over-diversity. To ensure correctness and consistency, demographic attribute assignments were fact-checked using reliable external sources such as Wikipedia and Wikidata. Demographic assignments fall into four categories: (i) examples with no assignment, where full diversity is expected; (ii) examples that inherently (linguistically) imply attributes (e.g., “actress” defined as a female actor/performer); (iii) institutional or historical rules (e.g., only baptized male persons can become pope); and (iv) examples referencing well-documented groups where missing attributes can be reasonably excluded (e.g., no female US presidents).

### v2

@@ Abstract
Current diversification strategies for text-to-image (T2I) models often ignore contextual appropriateness, leading to over-diversification where demographic attributes are modified even when explicitly specified in prompts. This paper introduces DivBench, a benchmark and evaluation framework for measuring both under- and over-diversification in T2I generation. Through systematic evaluation of state-of-the-art T2I models, we find that while most models exhibit limited diversity, many diversification approaches overcorrect by inappropriately altering contextually-specified attributes. We demonstrate that context-aware methods, particularly LLM-guided FairDiffusion and prompt rewriting, can already effectively address under-diversity while avoiding over-diversification, achieving a better balance between representation and semantic fidelity. [^1] We publicly release our data and code at anony.mous.
@@ 2.1 DivBench
We describe the overall setup in Fig. 1. We manually collected and annotated samples, divided into two main categories: (i) examples without assigned demographic attributes (e.g., “farmer” or “housekeeper”), which test whether models naturally reflect inclusive and diverse outputs, and (ii) examples with clearly defined demographic constraints (e.g., “actress” or “Maasai person”), which test whether models introduce inappropriate or historically inaccurate diversity, referred to as over-diversity. To ensure correctness and consistency, demographic attribute assignments were fact-checked using reliable external sources such as Wikipedia and Wikidata. We illustrate how to obtain such results with SPARQL from Wikidata in App. Fig. 4. Demographic assignments fall into four categories: (i) examples with no assignment, where full diversity is expected; (ii) examples that inherently (linguistically) imply attributes (e.g., “actress” defined as a female actor/performer); (iii) institutional or historical rules (e.g., only baptized male persons can become pope); and (iv) examples referencing well-documented groups where missing attributes can be reasonably excluded (e.g., no female US presidents).
@@ Appendix A Appendix
Determining Attributes Using Wikidata.
Fig. 4 presents example SQL/SPARQL queries used to retrieve data for specific groups from Wikidata. In Fig. 4(a), we illustrate how to query the gender property (wdt:P21) for the group ”King of France” (wd:Q842606). The query returns a list of all individuals categorized as Kings of France along with their associated gender, as shown in Fig. 4(b). In this case, all listed individuals are male, and therefore, we assign the protected attribute ”male” to this group.
@@
(a) Exemplary SQL/SPARQL query
(b) The first 15 results from the query output are shown; the full output is longer.
Figure 4: Wikidata queries with SQL/SPARQL and the obtained result. This example shows how to obtain information about attributes for specific groups.